OpenStudy (anonymous):

Use the elimination method to solve the system of equations. a +9b=21 a +4b=6 Is there one solution (if so simplify your answer) infinite solutions or no solutions?

OpenStudy (agent47):

a= 21-9b 21-9b+4b=6 21-5b=6 -5b=-15 b=3 a= 21-9*3 = 21-27 = -6

OpenStudy (agent47):

please tell me i got this thing right

OpenStudy (anonymous):

I don't know I am asking you for help LOL

OpenStudy (agent47):

that was directed @others

razor99 (razor99):

wait that is simulataneous equation right

OpenStudy (anonymous):

the answer is correct,,i've calculate it^^

OpenStudy (anonymous):

but i have to find an ordered pair.

OpenStudy (agent47):

(-6, 3)

OpenStudy (anonymous):

Thank you for showing me the steps and helping me Agent.